Assume that a procedure yields a binomial distribution with n trials and the probability of success for one trial is p.

n= 1575, p= 2/5

Use the range rule of thumb to find the minimum usual value μ−2σ and the maximum usual value μ+2σ.

μ = _____

σ = _____

μ−2σ = _____

μ+2σ = _____
Answer #1

GIVEN:

A procedure yields a binomial distribution with n trials and the probability of success for one trial is p.

n 1575

p = 215

FORMULAS USED:

In binomial distribution,

The formula for mean (mu ) is given by,

mu =np

The formula for standard deviation (sigma ) is given by,

sigma =[np(1-p)]^{1/2}

CALCULATION:

The mean (mu ) is given by,

mu =np

1575 * (2/5

μ 630

The standard deviation (sigma ) is given by,

sigma =[np(1-p)]^{1/2}

  1575 (2/5)(1 - (2/5)]

  630 0.6 /2

sigma =19.442

Thus the minimum usual value is given by,

μ-20 = 630-(2 * 19.442)

  =591.116591

The maximum usual value is given by,

mu +2sigma =630+(2*19.442)

  =668.884approx 669

Thus the minimum usual value is 591 and the maximum usual value is 669.